OpenStudy (anonymous):

What is the standard form of the equation of the line whose slope is -2 and which contains the point (3,-5)? show work

Nnesha (nnesha):

slope intercept form equation is y =mx + b first you have to find b do you know how ??

OpenStudy (anonymous):

y=2x+b Substitute -5 for y and 3 for x -5=2(3)+b -5=6+b -11=b y=-2x-11 is the slope intercept form

OpenStudy (anonymous):

no i hate slope questions

Nnesha (nnesha):

 plug this value and then solve for b